OpenStudy (anonymous):

Write an equation in point-slope form of the line through point J(4, 1) with slope –4.

OpenStudy (aripotta):

do you know what point-slope form is?

OpenStudy (anonymous):

yes

OpenStudy (aripotta):

so this should be easy. just substitute the numbers in

OpenStudy (anonymous):

how

OpenStudy (anonymous):

give u hard time

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

:(

OpenStudy (aripotta):

tell me what point-slope form is.

OpenStudy (anonymous):

tells a difference between two points on a line.

OpenStudy (aripotta):

what?

OpenStudy (anonymous):

The point-slope form expresses the fact that the difference in the y coordinate between two points on a line (that is, y − y1) is proportional to the difference in the x coordinate (that is, x − x1). The proportionality constant is m (the slope of the line).

OpenStudy (aripotta):

it's y - y1 = m(x - x1) idk if that's what you meant...

OpenStudy (aripotta):

brb i gotta walk mah dog

OpenStudy (anonymous):

at night

OpenStudy (anonymous):

im at USA

OpenStudy (aripotta):

alright i'm back

OpenStudy (aripotta):

so it should be easy. you're just substituting the y in the coordinate for y1, the x in the coordinate for x1, and -4 for m

OpenStudy (anonymous):

so the answer is y - y1 = m(x - x1)

OpenStudy (aripotta):

no, you didn't substitute

OpenStudy (aripotta):

our point is (4,1) so x1 is 4 and y1 is 1

OpenStudy (anonymous):

y - y1 = m(x - x1) y – 1 = (-4)(x – 4) y - 1 = -4(x - 4) y - 1 = -4x + 16 y = -4x + 16 + 1 y = -4x + 17

OpenStudy (aripotta):

which one is it then?

OpenStudy (anonymous):

y = -4x + 17

OpenStudy (anonymous):

STEp by STEp y - y1 = m(x - x1) y – 1 = (-4)(x – 4) y - 1 = -4(x - 4) y - 1 = -4x + 16 y = -4x + 16 + 1 y = -4x + 17

OpenStudy (aripotta):

i thought the answer was supposed to be in point-slope form

OpenStudy (anonymous):

so how can i write y = -4x + 17 to point-slope form

OpenStudy (aripotta):

lol. it was already in point-slope in step 1. y - 1 = -4(x - 4)

OpenStudy (anonymous):

haha never knew that